OpenStudy (anonymous):

simplify x^5 times x^3 times x^-1

OpenStudy (anonymous):

\[(x{}^{\wedge}5 )(x{}^{\wedge}3)\text{ }(x{}^{\wedge}-1)=\left(x^5 \right)\left(x^3\right)\text{ }\left(\frac{1}{x}\right)=x^7 \]